Hankins Corporation has 5.4 million shares of common stock outstanding; 290,000 shares of 5.2 percent preferred stock outstanding, par value of $100; and 125,000 5.7 percent semiannual bonds outstanding, par value $1,000 each. The common stock currently sells for $72 per share and has a beta of 1.13, the preferred stock currently sells for $103 per share, and the bonds have 20 years to maturity and sell for 103 percent of par. The market risk premium is 6.8 percent, T-bills are yielding 4.3 percent, and the firm’s tax rate is 23 percent. a. What is the firm's market value capital structure? (Do not round intermediate calculations and round your answers to 4 decimal places, e.g., .1616.) b. If the firm is evaluating a new investment project that has the same risk as the firm’s typical project, what rate should the firm use to discount the project’s cash flows? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1
Value Weight Cost
Equity 388.8 71% 11.98%
Preferred 29.87 5% 5.05%
Debt 128.75 24% 5.45%
Total 547.42 WACC 9.77%

Value of equity = 5.4 x 72 = 388.8 million

Value of preferred = 0.29 x 103 = 29.87 million

Vlaue of debt = 0.125 x 1030 = 128.75 million

Weight = Value / Total

Cost of equity, re = Rf + beta x MRP = 4.3% + 1.13 x 6.8% = 11.98%

Cost of preferred, rp = D / P = 5.2 / 103 = 5.05%

Cost of debt can be calculated using I/Y function on a calculator

N = 20 x 2 = 40, PMT = 5.7/2, PV = -103, FV = 100 => Compute I/Y = 2.73% (semi-annual)

Cost of debt, rd = 2.73% x 2 = 5.45%

WACC = wd x rd x (1 - tax) + wp x rp + we x re

= 24% x 5.45% x (1 - 23%) + 5% x 5.05% + 71% x 11.98%

= 9.77% is the discount rate